Telescope Focus Issues (hardware?)

Wednesday, February 5, 2020 6:56 AM

Problem(s) Encountered:

Focus got stuck again. Noticed it during evening twilight when I tried to run the focus4k prospero routine. No means of changing focus was able to pull it out. Shut everything down, including the TCS computer and chassis, and still no change. Previously, the focus would seem to come back on its own, but it did not this time. Noticed that there were some unusual noises in the dome. Eventually pinpointed it to the focus motor. It seemed to go away when I closed the TCS software, then start again when I reopened it and it was doing its temperature compensation. Decided to unplug the focus motor from the back of the chassis and observe at the last focus (5269). It varied from 5230-5310 which just happened to be fine for spectroscopy tonight. Lost about an hour.


With the telescope going off-sky for a few days, I decided to go ahead and pull the focus motor.  Even though things were working again for the remainder of this week, it seems worthwhile to try out the spare motor.  Wired it up and mounted it.  Testing shows it working as expected.  Justin has since tested it out.  After some initial complaining, the motor now appears to be working as expected (06Mar2020).